# Human as AI Mentor: Enhanced Human-in-the-loop Reinforcement Learning for Safe and Efficient Autonomous Driving

### v3

@@ 1 Introduction
Autonomous vehicles (AVs), with their potential to revolutionize the transportation industry, have garnered significant attention in recent years. This revolution promises enhanced road safety, optimized traffic flow, and improved fuel economy, as highlighted by recent studies (Feng et al., 2023; Dong et al., 2022; Shi et al., 2021, 2023; Jiang et al., 2022; Huang et al., 2023a; Han et al., 2023; Zhu et al., 2020; Wang et al., 2023). Yet, fully realizing these promises is fraught with challenges. Among these, one of the most critical is finding an appropriate driving policy that ensures AVs’ safety while optimizing traffic flow efficiency (Zhu et al., 2020; Zhu and Zhao, 2021; Di and Shi, 2021). In general, after undergoing extensive offline training, AVs are capable of generating their own online driving policies based on perception results from their sensors. These driving policies serve as the strategy or set of rules to which AVs adhere when navigating, including the operational (e.g., pedal and brake control, turn signals), tactical (e.g., lane-changing, lane-keeping), and strategic levels (e.g., routing) (Di and Shi, 2021). However, crafting a universally applicable driving policy for AVs, ensuring safe, efficient, and harmonious driving, is exceptionally challenging due to the dynamic and unpredictable nature of road environments. This complexity arises not only from the numerous possible traffic scenarios (Wu et al., 2023; Peng et al., 2022; Li et al., 2022c), but also from the intricate interactions between human-driven vehicles (HVs) and AVs on the roads (Shi et al., 2023; Chen et al., 2020).
@@ 1.1 IL and RL Methods for Driving Policy Learning
Driving policy learning is the main focus of this paper. Within the domain of learning-based driving policy learning, there are two main branches: Imitation Learning (IL) and Reinforcement Learning (RL) (Aradi, 2020). IL aims to emulate human driving behavior by replicating demonstrated actions in specific situations, such as behavior cloning (BC) (Le Mero et al., 2022), inverse RL (IRL) (Huang et al., 2023b), and generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016). During the demonstration phase, the novice agent avoids interacting with potentially risky environments, thereby ensuring training safety. However, two primary inherent problems are exposed: Firstly, distributional shift (Codevilla et al., 2019), where accumulated errors over time cause a deviation from the training distribution, leading to control failure. Secondly, the limitations of asymptotic performance (Silver et al., 2018), which are constrained by the proficiency of the demonstration source and unlikely to outperform it. In this context, conservative Q-learning (CQL) (Kumar et al., 2020) draws wide attention as it addresses the impact of out-of-distribution (OOD) actions in distributional shift problem. Conversely, RL employs a self-optimizing approach with the potential to address these imitation-related pitfalls (Aradi, 2020). RL-based methods, including proximal policy optimization (PPO) and soft actor-critic (SAC), have demonstrated successful applications across numerous challenging driving scenarios in AVs (Kiran et al., 2021).
In RL, human intentions are encoded into a reward function, enabling the agent to freely explore and learn from its environment. Moreover, reward-shaping (RS) methods, such as PPO-RS (Haarnoja et al., 2018) and SAC-RS (Schulman et al., 2017), have been proposed to address the issue of potentially diminished learning efficiency when the reward signals generated by the environment are sparse. However, safety during training and testing phase still presents a crucial challenge (Peng et al., 2022; Li et al., 2022c; Kiran et al., 2021; Muhammad et al., 2020). Due to the exploratory nature of RL, agents might inevitably encounter risky situations before learning how to avoid them. Constraint optimization-based methods have been developed to improve safety, such as trust region methods (e.g., constraint policy optimization, CPO (Achiam et al., 2017)) and Lagrangian methods (e.g., SAC-Lag (Ha et al., 2021), PPO-Lag (Stooke et al., 2020)). Nevertheless, these methods can only assure the upper bound of the agent’s failure probability, lacking an explicit mechanism to prevent the occurrence of critical failures (Li et al., 2022c). Moreover, the sampling efficiency of interaction between the agent and its environment is often low, leading to significant consumption of computational resources and extended training times (Wu et al., 2023).
@@ 1.2 Human-in-the-loop Learning Methods
Recently, some studies have attempted to incorporate human intelligence into the training loop of traditional RL or IL paradigms, leading to the development of the human-in-the-loop learning (HL) method (Christiano et al., 2017). This method, drawing upon the robustness and high adaptability of human beings in context understanding and knowledge-based reasoning, has been proven to enhance the safety and sampling efficiency of the conventional RL or IL training process. This integration has led to tangible successes in various real-world applications, such as robotic motion control (Krishna et al., 2022) and the development of large language models (Ouyang et al., 2022; Qu et al., 2023), with ChatGPT being a notable example. Generally, human intelligence in these contexts can manifest as human evaluation, human demonstration, and human intervention (Christiano et al., 2017; Wu et al., 2022, 2023; Li et al., 2022c). For example, in the robotics community, three representative methods of HL include dataset aggregation (DAgger) (Ross et al., 2011), human-gated DAgger (HG-DAgger) (Kelly et al., 2019), and intervention weighted regression (IWR) (Mandlekar et al., 2020). DAgger rectifies the compounding error of BC by intermittently requesting additional demonstrations from human experts. Instead of providing demonstrations upon request, HG-DAgger and IWR allow the human expert to intercede during exploration and guide the agent back to secure states.
While HL has made numerous achievements in the robotics community, its application in the realm of driving policy learning still faces several limitations: (a) Previous studies have mostly relied on human input in passive roles, such as suggesting which actions were favorable (Mandel et al., 2017) or evaluating collected trajectories according to human preference (Christiano et al., 2017). This passive human involvement can pose a risk to the human-AI collaborative system, as the agent explores the environment without sufficient safeguards. (b) Some studies train a human expert policy by utilizing offline human demonstrations (Nair et al., 2018). Nevertheless, the failure to make use of data from the agent’s free exploration can still result in OOD issues. (c) Many existing studies adhere to the traditional RL paradigm, which necessitates the design of a reward function (Ibarz et al., 2018). However, manually crafting a reward function that effectively encompasses all driving behaviors for a trained agent presents a significant challenge. (d) Although online takeover and demonstrations by humans could be more effective, long-term supervision can lead to exhaustion among human mentor (Wu et al., 2023; Li et al., 2022c). To accommodate the physical response of human drivers in the real-world, careful design must be taken on when and how human experts should be involved in the learning process.
@@
More importantly, as shown in Fig. 1 (a), many studies in the robotics community focus on the safety of individual AVs, such as collision avoidance under environmental uncertainties (Wu et al., 2022, 2023; Peng et al., 2022; Li et al., 2022c). Nonetheless, they often neglect the potential impact of these safety maneuvers employed for collision avoidance on subsequent traffic flow. This disregards a crucial factor in traffic management: traffic flow efficiency. On the other hand, many studies in the transportation community employ physics-based (rule-based) models to understand the impact of AVs on system-level performance, such as traffic flow disturbance (Chen et al., 2020; Eleonora et al., 2023; Ma et al., 2023), traffic congestion (Zhou et al., 2022; Du et al., 2023; Ha et al., 2023; Zhuo and Zhu, 2023; Yue et al., 2022), and traffic emission (Stern et al., 2019; Ansariyar and Tahmasebi, 2022). Meanwhile, some studies have begun integrating data-driven methodologies into rule-based models (Shi et al., 2023; Jiang et al., 2022; Han et al., 2023; Zhu et al., 2020; Chen et al., 2021b; Dong et al., 2021; Huang et al., 2018; Wu et al., 2019; Han et al., 2022; Ding et al., 2022; Liu et al., 2023; Dong et al., 2023). Although primarily targeting specific traffic scenarios, such as intersection optimization (Wu et al., 2019; Wu and Qu, 2022), ramp merging (Han et al., 2022; Zhu et al., 2022), and platoon control (Shi et al., 2023; Jiang et al., 2022), these models from the transportation community provide valuable insights into traffic management.
@@ 1.3 Contributions
Firstly, we propose a novel concept, the ‘$X+1+N$’ mixed traffic platoon, which integrate the dynamics of mixed traffic flow and environmental uncertainties, illustrated in Fig. 1 (a). The ‘$1+N$’ scenario, i.e., one leading AV, indexed as vehicle 0, followed by $N=\{1,\ldots,i,\ldots,N\}$ HVs, has already received extensive attention in the field of transportation (Shi et al., 2023; Jiang et al., 2022; Chen et al., 2021a; Mohammadian et al., 2023). The innovation in the paper is its consideration of environmental uncertainties, denoted ‘$X$’. This includes unforeseen elements in real-world traffic, such as alterations in road geometry (e.g., straight, ramp, roundabout, intersection), sudden appearance of other vehicles, and the presence of obstacles (e.g., traffic cones, warning triangles). These elements require the leading AV to react and maneuver instantly, such as accelerating or decelerating, to ensure its own safety. The actions taken by the AV in response to unforeseen elements ripple through to the following HVs, affecting the stability and efficiency of the entire platoon. Different from other studies in robotics community, we meticulously consider the impact of the AV’s actions on the following HVs, aiming to reduce traffic flow disturbance across the platoon.
Secondly, we reframe the role of humans within the learning process, transforming them from passive evaluators who merely provide preference assessments to active participants, as illustrated in Fig. 1 (b). In contrast to some studies that require human intervention by terminating episodes and resetting the environment, which may be impractical in the real world, we opt for a more intimate human-AI interaction approach, allowing humans and agents to share autonomy (Li et al., 2022c; Wu et al., 2023). Specifically, we integrate learning from intervention (LfI) and learning from demonstration (LfD) into a unified architecture. In this setup, human experts can intervene and override the actions of the original agents whenever they deem it necessary, and also provide real-time demonstrations through their actions. Moreover, the agent could be guided to avoid actions that cause significant traffic flow disturbance. Our key insight is that humans can serve as mentors for AI, and through the proposed HRL method, agents can rapidly learn to drive safely and efficiently in unseen scenarios.
@@ 2.2 Longitudinal Dynamical Modeling of Mixed Platoon
To model the dynamics of the ‘$X+1+N$’ mixed traffic platoon, we denote the position and velocity of HV $i$ at time $t$ as $loc_{\_}{i,t}^{HV}$, $v_{\_}{i,t}^{HV}$, respectively. The headway distance between vehicle $i$ and $i-1$, which is their relative (bumper-to-bumper) distance, is defined as $d_{\_}{i,t}^{HV}=loc_{\_}{i-1,t}^{HV}-loc_{\_}{i,t}^{HV}$. Additionally, $acc_{\_}{i,t}^{HV}$ represents the acceleration of vehicle $i$, and $\dot{d}_{\_}{i,t}^{HV}=v_{\_}{i-1,t}^{HV}-v_{\_}{i,t}^{HV}$ represents the relative velocity between HV $i$ and $i-1$. For simplicity, we disregard the vehicle length. Extensive research has meticulously dissected the car-following dynamics of HVs, resulting in the development of fundamental models such as the Optimal Velocity Model (OVM) (Bando et al., 1995) and the Intelligent Driver Model (IDM) (Treiber et al., 2000).
As demonstrated in Chen et al. (2021a), most of these models can be expressed in the following general form:
@@
To strike a balance between model fidelity and computational tractability, we adopt the IDM (Treiber et al., 2000) as the representative car-following model for the HVs in Fig. 1 (a). For the leading AV, it navigates using its throttle and steering signals, serving as the sole external inputs for the entire mixed platoon system.
@@ 3.2.1 Explicit Intervention Mechanism
The role of the explicit intervention mechanism is similar to that of an instructor in the driving school scenario. Specifically, human experts provide guidance, supervision, and direct intervention during the agent’s learning process. For this purpose, we adopt a technique termed the ‘switching function’ (Peng et al., 2022), which allows the agent to dynamically alternate between exploration and explicit intervention.
@@
Assumption 2 (Expert-level human takeover). While some studies attempt to establish an intervention threshold for identifying ’dangerous situations’ (Peng et al., 2022), this task is challenging due to the involvement of numerous non-quantifiable factors, as discussed by Wu et al. (2023). Drawing inspiration from the methods used by instructors to assess risky situations in driving school scenarios, we rely on human expert evaluations of safety during the training process rather than on external objective criteria. We assume that this expert-level human can accurately identify dangerous situations most of the time. In detail, the step-wise probability of that the human expert does not takeover when agents generates an unsafe action is bounded by a small value $\kappa<1$.
@@
Definition 1 (Probability-based Switch Function). The switch function $\mathcal{T}$ considers both the agent’s and human actions, thereby producing the modulated action $a_{\_}t^{mix}$, which is subsequently applied to the environment. Here. the intervention occurrence $\hat{c_{\_}t}$ indicates whether the human is in control. In this context, the intervention occurrence $\hat{c}$ indicates whether the human is in control. The function $\mathcal{T}(s_{\_}t,a_{\_}t,\pi_{\_}{human})$ determines which policy is responsible for the action, as discussed in Peng et al. (2022):

### v5

# HAIM-DRL: Enhanced Human-in-the-loop Reinforcement Learning for Safe and Efficient Autonomous Driving

@@ 1 Introduction
Autonomous vehicles (AVs), with their potential to revolutionize the transportation industry, have garnered significant attention in recent years. This revolution promises enhanced road safety, optimized traffic flow, and improved fuel economy, as highlighted by recent studies [Feng et al., 2023, Dong et al., 2022, Shi et al., 2021, 2023, Jiang et al., 2022, Huang et al., 2023a, Han et al., 2023, Zhu et al., 2020, Wang et al., 2023]. Yet, fully realizing these promises is fraught with challenges. Among these, one of the most critical is finding an appropriate driving policy that ensures AVs’ safety while optimizing traffic flow efficiency [Zhu et al., 2020, Zhu and Zhao, 2021, Di and Shi, 2021]. In general, after undergoing extensive offline training, AVs are capable of generating their own online driving policies based on perception results from their sensors. These driving policies serve as the strategy or set of rules to which AVs adhere when navigating, including the operational (e.g., pedal and brake control, turn signals), tactical (e.g., lane-changing, lane-keeping), and strategic levels (e.g., routing) [Di and Shi, 2021]. However, crafting a universally applicable driving policy for AVs, ensuring safe, efficient, and harmonious driving, is exceptionally challenging due to the dynamic and unpredictable nature of road environments. This complexity arises not only from the numerous possible traffic scenarios [Wu et al., 2023, Peng et al., 2022, Li et al., 2022c], but also from the intricate interactions between human-driven vehicles (HVs) and AVs on the roads [Shi et al., 2023, Chen et al., 2020].
@@ 1.1 IL and RL Methods for Driving Policy Learning
Driving policy learning is the main focus of this paper. Within the domain of learning-based driving policy learning, there are two main branches: Imitation Learning (IL) and Reinforcement Learning (RL) [Aradi, 2020]. IL aims to emulate human driving behavior by replicating demonstrated actions in specific situations, such as behavior cloning (BC) [Le Mero et al., 2022], inverse RL (IRL) [Huang et al., 2023b], and generative adversarial imitation learning (GAIL) [Ho and Ermon, 2016]. During the demonstration phase, the novice agent avoids interacting with potentially risky environments, thereby ensuring training safety. However, two primary inherent problems are exposed: Firstly, distributional shift [Codevilla et al., 2019], where accumulated errors over time cause a deviation from the training distribution, leading to control failure. Secondly, the limitations of asymptotic performance [Silver et al., 2018], which are constrained by the proficiency of the demonstration source and unlikely to outperform it. In this context, conservative Q-learning (CQL) [Kumar et al., 2020] draws wide attention as it addresses the impact of out-of-distribution (OOD) actions in distributional shift problem. Conversely, RL employs a self-optimizing approach with the potential to address these imitation-related pitfalls [Aradi, 2020]. RL-based methods, including proximal policy optimization (PPO) and soft actor-critic (SAC), have demonstrated successful applications across numerous challenging driving scenarios in AVs [Kiran et al., 2021].
In RL, human intentions are encoded into a reward function, enabling the agent to freely explore and learn from its environment. Moreover, reward-shaping (RS) methods, such as PPO-RS [Haarnoja et al., 2018] and SAC-RS [Schulman et al., 2017], have been proposed to address the issue of potentially diminished learning efficiency when the reward signals generated by the environment are sparse. However, safety during training and testing phase still presents a crucial challenge [Peng et al., 2022, Li et al., 2022c, Kiran et al., 2021, Muhammad et al., 2020]. Due to the exploratory nature of RL, agents might inevitably encounter risky situations before learning how to avoid them. Constraint optimization-based methods have been developed to improve safety, such as trust region methods (e.g., constraint policy optimization, CPO [Achiam et al., 2017]) and Lagrangian methods (e.g., SAC-Lag [Ha et al., 2021], PPO-Lag [Stooke et al., 2020]). Nevertheless, these methods can only assure the upper bound of the agent’s failure probability, lacking an explicit mechanism to prevent the occurrence of critical failures [Li et al., 2022c]. Moreover, the sampling efficiency of interaction between the agent and its environment is often low, leading to significant consumption of computational resources and extended training times [Wu et al., 2023].
@@ 1.2 Human-in-the-loop Learning Methods
Recently, some studies have attempted to incorporate human intelligence into the training loop of traditional RL or IL paradigms, leading to the development of the human-in-the-loop learning (HL) method [Christiano et al., 2017]. This method, drawing upon the robustness and high adaptability of human beings in context understanding and knowledge-based reasoning, has been proven to enhance the safety and sampling efficiency of the conventional RL or IL training process. This integration has led to tangible successes in various real-world applications, such as robotic motion control [Krishna et al., 2022] and the development of large language models [Ouyang et al., 2022, Qu et al., 2023], with ChatGPT being a notable example. Generally, human intelligence in these contexts can manifest as human evaluation, human demonstration, and human intervention [Christiano et al., 2017, Wu et al., 2022, 2023, Li et al., 2022c]. For example, in the robotics community, three representative methods of HL include dataset aggregation (DAgger) [Ross et al., 2011], human-gated DAgger (HG-DAgger) [Kelly et al., 2019], and intervention weighted regression (IWR) [Mandlekar et al., 2020]. DAgger rectifies the compounding error of BC by intermittently requesting additional demonstrations from human experts. Instead of providing demonstrations upon request, HG-DAgger and IWR allow the human expert to intercede during exploration and guide the agent back to secure states.
While HL has made numerous achievements in the robotics community, its application in the realm of driving policy learning still faces several limitations: (a) Previous studies have mostly relied on human input in passive roles, such as suggesting which actions were favorable [Mandel et al., 2017] or evaluating collected trajectories according to human preference [Christiano et al., 2017]. This passive human involvement can pose a risk to the human-AI collaborative system, as the agent explores the environment without sufficient safeguards. (b) Some studies train a human expert policy by utilizing offline human demonstrations [Nair et al., 2018]. Nevertheless, the failure to make use of data from the agent’s free exploration can still result in OOD issues. (c) Many existing studies adhere to the traditional RL paradigm, which necessitates the design of a reward function [Ibarz et al., 2018]. However, manually crafting a reward function that effectively encompasses all driving behaviors for a trained agent presents a significant challenge. (d) Although online takeover and demonstrations by humans could be more effective, long-term supervision can lead to exhaustion among human mentor [Wu et al., 2023, Li et al., 2022c]. To accommodate the physical response of human drivers in the real-world, careful design must be taken on when and how human experts should be involved in the learning process.
@@
More importantly, as shown in Fig. 1 (a), many studies in the robotics community focus on the safety of individual AVs, such as collision avoidance under environmental uncertainties [Wu et al., 2022, 2023, Peng et al., 2022, Li et al., 2022c]. Nonetheless, they often neglect the potential impact of these safety maneuvers employed for collision avoidance on subsequent traffic flow. This disregards a crucial factor in traffic management: traffic flow efficiency. On the other hand, many studies in the transportation community employ physics-based (rule-based) models to understand the impact of AVs on system-level performance, such as traffic flow disturbance [Chen et al., 2020, Eleonora et al., 2023, Ma et al., 2023], traffic congestion [Zhou et al., 2022, Du et al., 2023, Ha et al., 2023, Zhuo and Zhu, 2023, Yue et al., 2022], and traffic emission [Stern et al., 2019, Ansariyar and Tahmasebi, 2022]. Meanwhile, some studies have begun integrating data-driven methodologies into rule-based models [Shi et al., 2023, Jiang et al., 2022, Han et al., 2023, Zhu et al., 2020, Chen et al., 2021b, Dong et al., 2021, Huang et al., 2018, Wu et al., 2019, Han et al., 2022, Ding et al., 2022, Liu et al., 2023, Dong et al., 2023]. Although primarily targeting specific traffic scenarios, such as intersection optimization [Wu et al., 2019, Wu and Qu, 2022], ramp merging [Han et al., 2022, Zhu et al., 2022], and platoon control [Shi et al., 2023, Jiang et al., 2022], these models from the transportation community provide valuable insights into traffic management.
@@ 1.3 Contributions
Firstly, we propose a novel concept, the ‘$X+1+N$’ mixed traffic platoon, which integrate the dynamics of mixed traffic flow and environmental uncertainties, illustrated in Fig. 1 (a). The ‘$1+N$’ scenario, i.e., one leading AV, indexed as vehicle 0, followed by $N=\{1,\ldots,i,\ldots,N\}$ HVs, has already received extensive attention in the field of transportation [Shi et al., 2023, Jiang et al., 2022, Chen et al., 2021a, Mohammadian et al., 2023]. The innovation in the paper is its consideration of environmental uncertainties, denoted ‘$X$’. This includes unforeseen elements in real-world traffic, such as alterations in road geometry (e.g., straight, ramp, roundabout, intersection), sudden appearance of other vehicles, and the presence of obstacles (e.g., traffic cones, warning triangles). These elements require the leading AV to react and maneuver instantly, such as accelerating or decelerating, to ensure its own safety. The actions taken by the AV in response to unforeseen elements ripple through to the following HVs, affecting the stability and efficiency of the entire platoon. Different from other studies in robotics community, we meticulously consider the impact of the AV’s actions on the following HVs, aiming to reduce traffic flow disturbance across the platoon.
Secondly, we reframe the role of humans within the learning process, transforming them from passive evaluators who merely provide preference assessments to active participants, as illustrated in Fig. 1 (b). In contrast to some studies that require human intervention by terminating episodes and resetting the environment, which may be impractical in the real world, we opt for a more intimate human-AI interaction approach, allowing humans and agents to share autonomy [Li et al., 2022c, Wu et al., 2023]. Specifically, we integrate learning from intervention (LfI) and learning from demonstration (LfD) into a unified architecture. In this setup, human experts can intervene and override the actions of the original agents whenever they deem it necessary, and also provide real-time demonstrations through their actions. Moreover, the agent could be guided to avoid actions that cause significant traffic flow disturbance. Our key insight is that humans can serve as mentors for AI, and through the proposed HRL method, agents can rapidly learn to drive safely and efficiently in unseen scenarios.
@@ 2.2 Longitudinal Dynamical Modeling of Mixed Platoon
To model the dynamics of the ‘$X+1+N$’ mixed traffic platoon, we denote the position and velocity of HV $i$ at time $t$ as $loc_{\_}{i,t}^{HV}$, $v_{\_}{i,t}^{HV}$, respectively. The headway distance between vehicle $i$ and $i-1$, which is their relative (bumper-to-bumper) distance, is defined as $d_{\_}{i,t}^{HV}=loc_{\_}{i-1,t}^{HV}-loc_{\_}{i,t}^{HV}$. Additionally, $acc_{\_}{i,t}^{HV}$ represents the acceleration of vehicle $i$, and $\dot{d}_{\_}{i,t}^{HV}=v_{\_}{i-1,t}^{HV}-v_{\_}{i,t}^{HV}$ represents the relative velocity between HV $i$ and $i-1$. For simplicity, we disregard the vehicle length. Extensive research has meticulously dissected the car-following dynamics of HVs, resulting in the development of fundamental models such as the Optimal Velocity Model (OVM) [Bando et al., 1995] and the Intelligent Driver Model (IDM) [Treiber et al., 2000].
As demonstrated in Chen et al. [2021a], most of these models can be expressed in the following general form:
@@
To strike a balance between model fidelity and computational tractability, we adopt the IDM [Treiber et al., 2000] as the representative car-following model for the HVs in Fig. 1 (a). For the leading AV, it navigates using its throttle and steering signals, serving as the sole external inputs for the entire mixed platoon system.
@@ 3.2.1 Explicit Intervention Mechanism
The role of the explicit intervention mechanism is similar to that of an instructor in the driving school scenario. Specifically, human experts provide guidance, supervision, and direct intervention during the agent’s learning process. For this purpose, we adopt a technique termed the ‘switching function’ [Peng et al., 2022], which allows the agent to dynamically alternate between exploration and explicit intervention.
@@
Assumption 2 (Expert-level human takeover). While some studies attempt to establish an intervention threshold for identifying ’dangerous situations’ [Peng et al., 2022], this task is challenging due to the involvement of numerous non-quantifiable factors, as discussed by Wu et al. [2023]. Drawing inspiration from the methods used by instructors to assess risky situations in driving school scenarios, we rely on human expert evaluations of safety during the training process rather than on external objective criteria. We assume that this expert-level human can accurately identify dangerous situations most of the time. In detail, the step-wise probability of that the human expert does not takeover when agents generates an unsafe action is bounded by a small value $\kappa<1$.
@@
Definition 1 (Probability-based Switch Function). The switch function $\mathcal{T}$ considers both the agent’s and human actions, thereby producing the modulated action $a_{\_}t^{mix}$, which is subsequently applied to the environment. Here. the intervention occurrence $\hat{c_{\_}t}$ indicates whether the human is in control. In this context, the intervention occurrence $\hat{c}$ indicates whether the human is in control. The function $\mathcal{T}(s_{\_}t,a_{\_}t,\pi_{\_}{human})$ determines which policy is responsible for the action, as discussed in Peng et al. [2022]:
@@
$\mathcal{T}(s_{\_}t,a_{\_}t,\pi_{\_}{human})=\begin{cases}(a_{\_}t^{AV},1),&\text{if takeover }\\ \left(a_{\_}t^{human}\sim\pi_{\_}{human}(\cdot\mid s_{\_}t),0\right),&\text{ otherwise }\end{cases}$ (5)